OpenStudy (anonymous):

13 + w/7 = -18 and how did you do it?

OpenStudy (baldymcgee6):

13 + w/7 = -18 first move the 13 over to the other side. w/7 = -18 -13

OpenStudy (baldymcgee6):

then we multiply both sides by 7 to completely isolate w w = (-31)*7

OpenStudy (anonymous):

13 + w/7 = -18 first subtract 13 from both sides -13 -13 ---------------- w/7 = -31 Then multiply 7 by -31 = -217

OpenStudy (anonymous):

Thanks guys! :)